Granola is 2.10 per pound. Apples are 4.24 per pound. If you bought a 6 lb mix that cost 16.88, how many lbs of each did you buy?
Answer by Cromlix(4381)

x = granola y = apples
x + y = 6........1
210x + 424y = 1688.....2

Multiply 1 by 424
424x + 424y = 2544
210x + 424y = 1688
Subtract
214x = 856
x = 856/214
x = 4
Substitute x = 4 into 1
x + y = 6
4 + y = 6
y = 2
So 4 pounds of Granola and 2 pounds of apples